OpenStudy (anonymous):

What is the slope of the line that passes through the points (–1, 13) and (4, –2)?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large Slope = {\Delta y \over \Delta x} = {y_2 - y_1 \over x_2 - x_1} }\)

OpenStudy (anonymous):

m=y2-y1/x2-x1 m=-2-13/4+1

Parth (parthkohli):

Can you tell me what all co-ordinates \(x_1,x_2, y_1, y_2\) are?

OpenStudy (anonymous):

m=-15/5, m=-3

OpenStudy (anonymous):

\[\huge \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

It's simply plugging in the numbers into the formula...:).

OpenStudy (anonymous):

thanks! i actually get it more than i did before

OpenStudy (anonymous):

That's great!!! And you're welcome! :).

OpenStudy (anonymous):

And as you start off you can label your points y2, y1, x2, x1 until you can do it out of your head :).